# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
(cf. [72], Corollary A.2 in [31], and Corollary 9.3 in [20])   Given a small positive constant $\varepsilon$, there exist positive constants $C_{1}^{\prime}=C_{1}^{\prime}(n,\varepsilon)$, $C_{2}^{\prime}=C_{2}^{\prime}(n,\varepsilon)$ and $C_{3}^{\prime}=C_{3}^{\prime}(n)$ with the following property: For each space-time point $(x_{0},t_{0})$ on an orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$ (with at most an isolated singularity), there is an open subset $U$ of $\mathbb{R}^{n}//\Gamma$ with $B(x_{0},t_{0},C_{1}^{\prime-1}R(x_{0},t_{0})^{-\frac{1}{2}})\subset U\subset B(x_{0},t_{0},C_{1}^{\prime}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{\prime-1}R(x_{0},t_{0})$ and $C_{2}^{\prime}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$, and on $U\times[t_{0}-\min\{R(x_{0},t_{0})^{-1},t_{0}\},t_{0}]$ the solution $\hat{g}_{\Gamma}(t)$ is, after scaling with $R(x_{0},t_{0})$ and shifting $t_{0}$ to zero, $\varepsilon$-close to the corresponding subset of the evolving round cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ over the time interval $[-\min\{t_{0}R(x_{0},t_{0}),1\},0]$ with scalar curvature 1 at the time zero, and $U$ is disjoint from the surgery cap (in particular, $x_{0}\notin B(p_{\Gamma},0,\varepsilon^{-1})$) when $t_{0}\leq R(x_{0},t_{0})^{-1}$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ (diffeomorphic to $\mathbb{R}^{n}//\Gamma$);
@@
Proof.    The conclusion in the case that $\Gamma$ is trivial follows from Corollary A.2 in [31], Corollary 9.3 in [20], and Proposition 2.6. Now we consider the case that $\Gamma$ is not trivial. By Corollary A.2 in [31] and Corollary 9.3 in [20], for any $\eta>0$ there exist positive constants $\tilde{C}_{1}=\tilde{C}_{1}(n,\eta)$, $\tilde{C}_{2}=\tilde{C}_{2}(n,\eta)$ and $C_{3}^{\prime}=C_{3}^{\prime}(n)$ with the following property: For each space-time point $(\tilde{x}_{0},t_{0})$ on the standard solution $(\mathbb{R}^{n},\hat{g}(t))$, there is an open subset $\tilde{U}$ of $\mathbb{R}^{n}$ with $B(\tilde{x}_{0},t_{0},\tilde{C}_{1}^{-1}R(\tilde{x}_{0},t_{0})^{-\frac{1}{2}})\subset\tilde{U}\subset B(\tilde{x}_{0},t_{0},\tilde{C}_{1}R(\tilde{x}_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $\tilde{U}$ at time $t_{0}$ lying between $\tilde{C}_{2}^{-1}R(\tilde{x}_{0},t_{0})$ and $\tilde{C}_{2}R(\tilde{x}_{0},t_{0})$, which falls into one of the following two categories:
@@
($\tilde{b}$) $\tilde{U}$ is an $\eta$-cap centered at $(\tilde{x}_{0},t_{0})$;
moreover, the scalar curvature in $\tilde{U}$ at time $t_{0}$ satisfies the derivative estimates
@@
Now by above and Lemma C2.1, Corollary C2.3 and Theorem C2.4 in [48] one can get a Hamilton’s canonical parametrization
@@
where $C(n,\eta)$ is a suitable positive constant. As $(\mathbb{R}^{n},\hat{g}(t_{0}))$ is $O(n)$-invariant, $\tilde{\phi}$ is $\Gamma$-equivariant, where the $\Gamma$-action on the second factor of $\mathbb{S}^{n-1}\times(a,+\infty)$ is trivial. If $\eta$ is sufficiently small (depending on $\varepsilon$ and $n$), the quotient of $\tilde{\phi}$ by $\Gamma$, denoted by
@@
will give a $\beta\varepsilon$-tube, where $\beta$ is from Lemma 5.8. (Compare Corollary C2.3 in [48] and the proof of Lemma 5.6.) Moreover, if $(x_{0},t_{0})$ is a center of a $\beta\varepsilon$-neck contained in this $\beta\varepsilon$-tube with $t_{0}>\frac{n}{n+2}$, which implies that $t_{0}R(x_{0},t_{0})\geq t_{0}\frac{1}{1-\frac{2}{n}t_{0}}>1$, by using Lemma 5.8 (choosing $b=0$ there) we see that $(x_{0},t_{0})$ will fall into the category (a); if $(x_{0},t_{0})$ is a center of a $\beta\varepsilon$-neck contained in this $\beta\varepsilon$-tube with $t_{0}\leq\frac{n}{n+2}$, by using the fact that the orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(\cdot))$ restricted to the time interval $[0,\frac{n}{n+2}]$ is asymptotic to the evolving round quotient cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ (with scalar curvature 1 at time 0) restricted to the same time interval at infinity, which can be shown, for example, by working $O(n)$-equivariantly with a manifold standard solution and using the arguments in the last paragraph on p. 262 of [31] and in Section 3.2 of [26], we see that $(x_{0},t_{0})$ will still fall into the category (a).
@@ 3 Existence of $(r,\delta)$-surgical solutions
The following lemma should be known (for relevant results in the manifold case see for example Proposition 7.7 in [3], Theorem 2.2 in Chapter 8 of [49], and p. 76 of [29]).
@@
Then $f_{3}$ is a homeomorphism. Using the ambient tubular neighborhood theorem (actually here it suffices to apply the ambient tubular neighborhood theorem to the manifold $P\cup_{\psi}\partial_{1}P\times[0,1]\setminus\{\text{singularities}\}$) and arguing as in the proof of Theorem 1.9 in Chapter 8 of [49], we see that $f_{3}$ can be modified to a diffeomorphism $f_{4}:P\rightarrow P\cup_{\psi}\partial_{1}P\times[0,1]$. $\Box$
@@
Let $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7, and the constant $\varepsilon_{4}$ is as in Proposition 2.5. Fix $0<\varepsilon<\tilde{\varepsilon}_{1}(\varepsilon_{0})$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6. Let $(\mathcal{O},g)$ be a closed, connected $n$-orbifold with at most isolated singularities. If each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck or an $\varepsilon$-cap, then $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a weak connected sum of at most two spherical orbifolds with at most isolated singularities. In particular, if each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck, then $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries.
Proof. Compare the proof of Proposition 8.1 in [20]. Let $\mathcal{O}$ satisfy the assumption of our proposition. If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{R}^{n}$ but no $\varepsilon$-caps of the other types, by using the definition of $\varepsilon$-cap of type $\mathbb{R}^{n}$ (and our choice of $\varepsilon$), Hamilton’s canonical parametrization of an $\varepsilon$-tube (cf. Section C.2 in [48]; see also Lemma 5.6 and Proposition 5.7 in the appendix), and Theorem 1.5 in Chapter 8 of [49] (a corollary of the isotopy extension theorem), we see that it is diffeomorphic to $D^{n}\cup_{h}D^{n}$, where $h$ is an isometry of $\mathbb{S}^{n-1}$, so it is diffeomorphic to $\mathbb{S}^{n}$. If $\mathcal{O}$ contains an $\varepsilon$-cap of type $\mathbb{R}^{n}$ and an $\varepsilon$-cap diffeomorphic to $\mathbb{R}P^{n}\setminus\bar{B}$, by using the definition of these two $\varepsilon$-caps, Hamilton’s canonical parametrization of an $\varepsilon$-tube, and Theorem 1.5 in Chapter 8 of [49], we see that it is diffeomorphic to $(\mathbb{R}P^{n}\setminus B)\cup_{h}D^{n}$, where $h:\partial(\mathbb{R}P^{n}\setminus B)\rightarrow\partial D^{n}$ is a homothety, so it is diffeomorphic to $\mathbb{R}P^{n}$. Similarly, if $\mathcal{O}$ contains a smooth $\varepsilon$-cap of type $\mathbb{R}^{n}$ and an orbifold $\varepsilon$-cap of type II, it is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})$.
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $C_{\Gamma}$ ($|\Gamma|\geq 2$) but no $\varepsilon$-caps of the other types, by using Proposition 5.4 and Lemma 3.1 we see that it is diffeomorphic to $D^{n}//\Gamma\cup_{h}D^{n}//\Gamma$, where $h$ is a diffeomorphism of $\mathbb{S}^{n-1}/\Gamma$, and can be viewed as a weak connected sum of the form $\mathbb{S}^{n}//\Gamma\tilde{\sharp}\mathbb{S}^{n}//\Gamma$ (where the weak connected sum occurs at two singular points).
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $C_{\Gamma}^{\sigma}$ ($|\Gamma|\geq 2$) but no caps of the other types, by using Proposition 5.4 and Lemma 3.1 we see that it is a smooth manifold diffeomorphic to $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B)\cup_{h}(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime})$ (the notation is as in Section 2), where $h$ is a diffeomorphism between the boundaries, and has the structure of an orbifiber bundle over $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$. Clearly it can be viewed as a weak connected sum of the form $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\tilde{\sharp}\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle$ (where the weak connected sum occurs at two singular points).
By using Lemma 5.6, Proposition 5.7 in the appendix and Theorem C2.5 in [48] we see that if each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck, then $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries, and has the structure of a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$ with structure group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ (by the way, when the fiber is diffeomorphic to $\mathbb{S}^{n-1}$ or $\mathbb{R}P^{n-1}$, one can also consult [42] or [38]); in this case it can be viewed as a connected sum on a single orbifold $\mathbb{S}^{n}//\Gamma$, that is, a connected sum of the form $\mathbb{S}^{n}//\Gamma\sharp_{f}$. $\Box$
@@
[…]
@@
Pinching assumption (see Definition 7.1 in [20]): Let $f:[0,\infty)\rightarrow[0,\infty)$ be an increasing concave function satisfying $\text{lim}_{s\rightarrow\infty}\frac{f(s)}{s}=0$, and $\theta$ be a positive number. A Riemannian orbifold satisfies the $(f,\theta)$-pinching assumption (or has $(f,\theta)$-pinched curvature) if $\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in$ PIC2 and $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in$ PIC. An evolving Riemannian orbifold $(\mathcal{O}(t),g(t))$, $t\in I$, has $(f,\theta)$-pinched curvature if it has $(f,\theta)$-pinched curvature at each time $t\in I$.
@@
Definition (cf. [72] and p. 841 in [28]). Let $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$ be as in the definition of $\varepsilon$-caps in Section 3, and $C$ be a positive constant. A point $(x,t)$ in a surgical solution to the Ricci flow is said to have an $(\varepsilon,C)$-canonical neighborhood if either the point $x$ is contained in a compact component which is strictly PIC2 at time $t$, or there is an open neighborhood $U$ of $x$ satisfying $B(x,t,C^{-1}R(x,t)^{-\frac{1}{2}})\subset U\subset B(x,t,CR(x,t)^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t$ lying (strictly) between $C^{-1}R(x,t)$ and $CR(x,t)$, which falls into one of the following two types:
(a) $U$ is a strong $\varepsilon$-neck centered at $(x,t)$,
(b) $U$ is an $\varepsilon$-cap centered at $(x,t)$,
and moreover, the scalar curvature in $U$ at time $t$ satisfies the derivative estimates

### v22

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
(cf. [76], Corollary A.2 in [33], and Corollary 9.3 in [21])   Given a small positive constant $\varepsilon$, there exist positive constants $C_{1}^{\prime}=C_{1}^{\prime}(n,\varepsilon)$, $C_{2}^{\prime}=C_{2}^{\prime}(n,\varepsilon)$ and $C_{3}^{\prime}=C_{3}^{\prime}(n)$ with the following property: For each space-time point $(x_{0},t_{0})$ on an orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$ (with at most an isolated singularity), there is an open subset $U$ of $\mathbb{R}^{n}//\Gamma$ with $\overline{B(x_{0},t_{0},C_{1}^{\prime-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}^{\prime}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{\prime-1}R(x_{0},t_{0})$ and $C_{2}^{\prime}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is an $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, and on $U\times[t_{0}-\min\{R(x_{0},t_{0})^{-1},t_{0}\},t_{0}]$ the solution $\hat{g}_{\Gamma}(t)$ is, after scaling with $R(x_{0},t_{0})$ and shifting $t_{0}$ to zero, $\varepsilon$-close to the corresponding subset of the evolving round cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ over the time interval $[-\min\{t_{0}R(x_{0},t_{0}),1\},0]$ with scalar curvature 1 at the time zero, and $U$ is disjoint from the surgery cap (in particular, $x_{0}\notin B(p_{\Gamma},0,\varepsilon^{-1})$, where, as in the appendix, $p_{\Gamma}$ denotes the tip of the orbifold standard solution) when $t_{0}\leq R(x_{0},t_{0})^{-1}$, or
(b) $U$ is an $\varepsilon$-cap diffeomorphic to $\mathbb{R}^{n}//\Gamma$ centered at $(x_{0},t_{0})$,
and $vol_{\hat{g}_{\Gamma}(t_{0})}(U)>(C_{2}^{\prime}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2}$; moreover, the scalar curvature in $U$ at time $t_{0}$ satisfies the derivative estimates
@@
Proof.    The conclusion in the case that $\Gamma$ is trivial follows from Corollary A.2 in [33], Corollary 9.3 in [21], and Proposition 2.6. Now we consider the case that $\Gamma$ is not trivial. By Corollary A.2 in [33] and Corollary 9.3 in [21], for any $\eta>0$ there exist positive constants $\tilde{C}_{1}=\tilde{C}_{1}(n,\eta)$, $\tilde{C}_{2}=\tilde{C}_{2}(n,\eta)$ and $C_{3}^{\prime}=C_{3}^{\prime}(n)$ with the following property: For each space-time point $(\tilde{x}_{0},t_{0})$ on the standard solution $(\mathbb{R}^{n},\hat{g}(t))$, there is an open subset $\tilde{U}$ of $\mathbb{R}^{n}$ with $\overline{B(\tilde{x}_{0},t_{0},\tilde{C}_{1}^{-1}R(\tilde{x}_{0},t_{0})^{-\frac{1}{2}})}\subset\tilde{U}\subset B(\tilde{x}_{0},t_{0},\tilde{C}_{1}R(\tilde{x}_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $\tilde{U}$ at time $t_{0}$ lying between $\tilde{C}_{2}^{-1}R(\tilde{x}_{0},t_{0})$ and $\tilde{C}_{2}R(\tilde{x}_{0},t_{0})$, which falls into one of the following two categories:
@@
($\tilde{b}$) $\tilde{U}$ is an $\eta$-cap centered at $(\tilde{x}_{0},t_{0})$,
and $\text{vol}_{\hat{g}(t_{0})}(\tilde{U})>\tilde{C}_{2}^{-1}R(\tilde{x}_{0},t_{0})^{-n/2}$; moreover, the scalar curvature in $\tilde{U}$ at time $t_{0}$ satisfies the derivative estimates
@@
Now by above and Lemma C2.1, Corollary C2.3 and Theorem C2.4 in [51] (see also the appendix) one can get a Hamilton’s canonical parametrization
@@
where $p_{0}$ denotes the tip of the manifold standard solution as in the appendix, and $C(n,\eta)$ is a large number. As $(\mathbb{R}^{n},\hat{g}(t_{0}))$ is $O(n)$-invariant, $\tilde{\phi}$ is $\Gamma$-equivariant, where the $\Gamma$-action on the second factor of $\mathbb{S}^{n-1}\times(a,+\infty)$ is trivial. If $\eta$ is sufficiently small (depending on $\varepsilon$ and $n$), the quotient of $\tilde{\phi}$ by $\Gamma$, denoted by
@@
will give a $\beta\varepsilon$-tube, where $\beta$ is from Lemma A.8. (Compare Corollary C2.3 in [51] and the proof of Lemma A.6.) Moreover, if $(x_{0},t_{0})$ is a center of a $\beta\varepsilon$-neck contained in this $\beta\varepsilon$-tube with $t_{0}>\frac{n}{n+2}$, which implies that $t_{0}R(x_{0},t_{0})\geq t_{0}\frac{1}{1-\frac{2}{n}t_{0}}>1$, by using Lemma A.8 (choosing $b=0$ there) we see that $(x_{0},t_{0})$ will fall into the category (a); if the constant $C(n,\eta)$ above is sufficiently large (depending on $\varepsilon$) and $(x_{0},t_{0})$ is a center of a $\beta\varepsilon$-neck contained in this $\beta\varepsilon$-tube with $t_{0}\leq\frac{n}{n+2}$, by using the fact that the orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(\cdot))$ restricted to the time interval $[0,\frac{n}{n+2}]$ is asymptotic to the evolving round quotient cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ (with scalar curvature 1 at time 0) restricted to the same time interval at infinity, which can be shown, for example, by working $O(n)$-equivariantly with a manifold standard solution and using the arguments in the last paragraph on p. 262 of [33] and in Section 3.2 of [28], we see that $(x_{0},t_{0})$ will still fall into the category (a).
@@ 3 Existence of $(r,\delta)$-surgical solutions
The following lemma should be known (for relevant results in the manifold case see for example Proposition 7.7 in [3], Theorem 2.2 in Chapter 8 of [52], Proposition 7.6.6 in [71] and p. 76 of [31]).
@@
Then $f_{3}$ is a homeomorphism. Using the ambient tubular neighborhood theorem (actually here it suffices to apply the ambient tubular neighborhood theorem to the manifold $P\cup_{\psi}\partial_{1}P\times[0,1]\setminus\{\text{singularities}\}$) and arguing as in the proof of Theorem 1.9 in Chapter 8 of [52] and Proposition 7.6.4 in [71], we see that $f_{3}$ can be modified to a diffeomorphism $f_{4}:P\rightarrow P\cup_{\psi}\partial_{1}P\times[0,1]$. $\Box$
@@
Let $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition A.7, and the constant $\varepsilon_{4}$ is as in Proposition 2.5. Fix $0<\varepsilon<\tilde{\varepsilon}_{1}(\varepsilon_{0})$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma A.6. Let $(\mathcal{O},g)$ be a closed, connected $n$-orbifold with at most isolated singularities. If each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck or an $\varepsilon$-cap, then $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a weak connected sum of at most two spherical orbifolds with at most isolated singularities. In particular, if each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck, then $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries.
Proof. Compare the proof of Proposition 8.1 in [21]. Let $\mathcal{O}$ satisfy the assumption of our proposition. If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{R}^{n}$ but no $\varepsilon$-caps of the other types, by using the definition of $\varepsilon$-cap of type $\mathbb{R}^{n}$ (and our choice of $\varepsilon$), Hamilton’s canonical parametrization of an $\varepsilon$-tube (cf. Section C.2 in [51]; see also Lemma A.6 and Proposition A.7 in the appendix), and Theorem 1.5 in Chapter 8 of [52] (a corollary of the isotopy extension theorem), we see that it is diffeomorphic to $D^{n}\cup_{h}D^{n}$, where $h$ is an isometry of $\mathbb{S}^{n-1}$, so it is diffeomorphic to $\mathbb{S}^{n}$. (To be more precise, here we also need to use, for example, a connectedness argument. Let $W$ be a topological cap which is the union of an $\varepsilon$-cap and an $\varepsilon_{0}$-tube (with Hamilton’s canonical parametrization). Then a cross section $S$ of the $\varepsilon_{0}$-tube is separating in $W$, and it divides $W$ into two (connected) parts, both with boundary $S$. Let $X$ be an embedded compact, connected suborbifold (of codimension 0) in $W$ with boundary $S$. Then $X\setminus S$ is still path-connected. It follows that $X$ is contained in one of the two parts of $W$ divided by $S$. Since $X$ is compact (and closed), and open relative to the part it lies in, it must occupy a whole part (of course it is the compact part).) If $\mathcal{O}$ contains an $\varepsilon$-cap of type $\mathbb{R}^{n}$ and an $\varepsilon$-cap diffeomorphic to $\mathbb{R}P^{n}\setminus\bar{B}$, by using the definition of these two $\varepsilon$-caps A cross section of Hamilton’s canonical parametrization should separate the union of an $\varepsilon$-cap and an $\varepsilon_{0}$-tube with Hamilton’s canonical parametrization. If $\mathcal{O}$ contains an $\varepsilon$-cap of type $\mathbb{R}^{n}$ and an $\varepsilon$-cap diffeomorphic to $\mathbb{R}P^{n}\setminus\bar{B}$, by using the definition of these two $\varepsilon$-caps, Hamilton’s canonical parametrization of an $\varepsilon$-tube, and Theorem 1.5 in Chapter 8 of [52], we see that it is diffeomorphic to $(\mathbb{R}P^{n}\setminus B)\cup_{h}D^{n}$, where $h:\partial(\mathbb{R}P^{n}\setminus B)\rightarrow\partial D^{n}$ is a homothety, so it is diffeomorphic to $\mathbb{R}P^{n}$. Similarly, if $\mathcal{O}$ contains a smooth $\varepsilon$-cap of type $\mathbb{R}^{n}$ and an orbifold $\varepsilon$-cap of type II, it is diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})$.
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $C_{\Gamma}$ ($|\Gamma|\geq 2$) but no $\varepsilon$-caps of the other types, by using Proposition A.4 and Lemma 3.1 we see that it is diffeomorphic to $D^{n}//\Gamma\cup_{h}D^{n}//\Gamma$, where $h$ is a diffeomorphism of $\mathbb{S}^{n-1}/\Gamma$, and can be viewed as a weak connected sum of the form $\mathbb{S}^{n}//\Gamma\tilde{\sharp}\mathbb{S}^{n}//\Gamma$ (where the weak connected sum occurs at two singular points).
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $C_{\Gamma}^{\sigma}$ ($|\Gamma|\geq 2$) but no caps of the other types, by using Proposition A.4 and Lemma 3.1 we see that it is a smooth manifold diffeomorphic to $(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B)\cup_{h}(\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle\setminus B^{\prime})$ (the notation is as in Section 2), where $h$ is a diffeomorphism between the boundaries, and has the structure of an orbifiber bundle over $\mathcal{I}$ with generic fiber diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$. Clearly it can be viewed as a weak connected sum of the form $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\tilde{\sharp}\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma^{\prime}}\rangle$ (where the weak connected sum occurs at two singular points).
By using Lemma A.6, Proposition A.7 in the appendix and Theorem C2.5 in [51] we see that if each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck, then $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries, and has the structure of a $\mathbb{S}^{n-1}/\Gamma$-bundle over $\mathbb{S}^{1}$ with structure group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$ (by the way, when the fiber is diffeomorphic to $\mathbb{S}^{n-1}$ or $\mathbb{R}P^{n-1}$, one can also consult [45] or [41]); in this case it can be viewed as a connected sum on a single orbifold $\mathbb{S}^{n}//\Gamma$, that is, a connected sum of the form $\mathbb{S}^{n}//\Gamma\sharp_{f}$. $\Box$
@@
$\mathcal{F}_{t}\subset\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}\}\cap\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}2}\}$ (3.1)
@@
Pinching assumption (see Definition 7.1 in [21]): Let $f:[0,\infty)\rightarrow[0,\infty)$ be an increasing concave function satisfying $\text{lim}_{s\rightarrow\infty}\frac{f(s)}{s}=0$, and $\theta$ be a positive number. A Riemannian orbifold satisfies the $(f,\theta)$-pinching assumption (or has $(f,\theta)$-pinched curvature) if $\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}2}$ and $\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$. An evolving Riemannian orbifold $(\mathcal{O}(t),g(t))$, $t\in I$, has $(f,\theta)$-pinched curvature if it has $(f,\theta)$-pinched curvature at each time $t\in I$.
@@
Definition (cf. [76] and p. 841 in [30]). Let $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$ be as in the definition of $\varepsilon$-caps in Section 2, and $C$ be a positive constant. A space-time point point $(x,t)$ in a surgical solution $g(\cdot)$ to the $n$-dimensional Ricci flow is said to have an $(\varepsilon,C)$-canonical neighborhood if either the point $x$ is contained in a compact component which is strictly PIC2 at time $t$, or there is an open neighborhood $U$ of $x$ satisfying $\overline{B(x,t,C^{-1}R(x,t)^{-\frac{1}{2}})}\subset U\subset B(x,t,CR(x,t)^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t$ lying (strictly) between $C^{-1}R(x,t)$ and $CR(x,t)$, which falls into one of the following two types:
(a) $U$ is a strong $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x,t)$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x,t)$ with any cross-section of the $\varepsilon$-neck contained in it diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$
for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and
$\text{vol}_{g(t)}(U)>(C|\Gamma|)^{-1}R(x,t)^{-n/2};$
moreover, the scalar curvature in $U$ at time $t$ satisfies the derivative estimates